### v2

@@ 4.2 New left and right preconditioner
$~{}~{}~{}~{}\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$
$=\frac{\max\{\underset{i,j\in[m],i\neq j}{\max}\bar{v}_{ij}(\delta),\frac{\mu_{1}(\sigma_{1}+\sigma_{m+1})}{\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}}\}}{\min\{\underset{i,j\in[m],i\neq j}{\min}\underline{v}_{ij}(\delta),\frac{\mu_{m}(\sigma_{m}-\sigma_{m+1})}{\sqrt{\mu_{m}^{2}\sigma_{m}^{2}+\delta}}\}}$
$=\frac{\max\{\bar{v}_{12}(\delta),\frac{\mu_{1}(\sigma_{1}+\sigma_{m+1})}{\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}}\}}{\min\{\underline{v}_{12}(\delta),\frac{\mu_{m}(\sigma_{m}-\sigma_{m+1})}{\sqrt{\mu_{m}^{2}\sigma_{m}^{2}+\delta}}\}}$
(11) $=\frac{\max\{\frac{1}{2}(\mu_{1}+\mu_{2})(\sigma_{1}+\sigma_{2}),\frac{\mu_{1}(\sigma_{1}+\sigma_{m+1})}{2\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}}(\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}+\sqrt{\mu_{2}^{2}\sigma_{2}^{2}+\delta})\}}{\min\{\frac{1}{2}(\mu_{1}-\mu_{2})(\sigma_{1}-\sigma_{2}),\frac{\mu_{m}(\sigma_{m}-\sigma_{m+1})}{2\sqrt{\mu_{m}^{2}\sigma_{m}^{2}+\delta}}(\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}+\sqrt{\mu_{2}^{2}\sigma_{2}^{2}+\delta})\}}$
(12) $\leq\frac{\max\{\frac{1}{2}(\mu_{1}+\mu_{2})(\sigma_{1}+\sigma_{2}),\mu_{1}(\sigma_{1}+\sigma_{m+1})\}}{\min\{\underset{i,j\in[m],i\neq j}{\min}\frac{1}{2}(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j}),\mu_{m}(\sigma_{m}-\sigma_{m+1})\}}$
@@
$=\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*})),$
where we multiply both the numerator and the denominator by $\frac{1}{2}(\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}+\sqrt{\mu_{2}^{2}\sigma_{2}^{2}+\delta})$ to yield Eq. 11, and we use the fact that $2\sqrt{\mu_{m}^{2}\sigma_{m}^{2}+\delta}<\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}+\sqrt{\mu_{2}^{2}\sigma_{2}^{2}+\delta}<2\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}$ to yield Eq. 12.
@@ 4.3 RGD and RCG for canonical correlation analysis
Note that 1) the retraction mapping is the generalized QR decomposition [22] with respect to $\Sigma_{xx}$ and $\Sigma_{yy}$, i.e.,
@@
where $\operatorname{qf}(\mathbf{X})$ refers to the $\mathbf{Q}$ factor in the QR decomposition $\mathbf{Q}\mathbf{R}=\mathbf{X}$. In practice, the retraction can be efficiently computed [22] by $\operatorname{R}_{(\mathbf{U},\mathbf{V})}(\eta)=((\mathbf{U}+\eta_{1})\mathbf{R}_{1}^{-1},(\mathbf{V}+\eta_{2})\mathbf{R}_{2}^{-1})$ instead, where $\mathbf{R}_{1}^{\mathsf{T}}\mathbf{R}_{1}=(\mathbf{U}+\eta_{1})^{\mathsf{T}}\Sigma_{xx}(\mathbf{U}+\eta_{1})$ and $\mathbf{R}_{2}^{\mathsf{T}}\mathbf{R}_{2}=(\mathbf{V}+\eta_{2})^{\mathsf{T}}\Sigma_{yy}(\mathbf{V}+\eta_{2})$ are Cholesky decompositions; 2) the vector transport in Algorithm 4 is defined by the projection operator in Eq. 8, i.e., $\mathcal{T}_{t\leftarrow t-1}(\eta)=\Pi_{g,(\mathbf{U}^{(t)},\mathbf{V}^{(t)})}(\eta)$ for $\eta\in\operatorname{T}_{(\mathbf{U}^{(t-1)},\mathbf{V}^{(t-1)})}\!\mathcal{M}$.
@@ 4.4 Numerical validation
5 Appplication to truncated singular value decomposition
@@ 5.1 A new preconditioned metric
The effect of the new metric Eq. 2 is illustrated by the following proposition, which can be proved in a same fashion as Propositions 4.1, 4.4, and 4.8.
@@
$\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ $=\frac{\max\{\underset{i,j\in[p],i\neq j}{\max}\frac{(\mu_{i}+\mu_{j})(\sigma_{i}+\sigma_{j})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta}},\frac{\mu_{1}(\sigma_{1}+\sigma_{p+1})}{\sqrt{\sigma_{1}^{2}\mu_{1}^{2}+\delta}}\}}{\min\{\underset{i,j\in[p],i\neq j}{\min}\frac{(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta}},\frac{\mu_{p}(\sigma_{p}-\sigma_{p+1})}{\sqrt{\sigma_{p}^{2}\mu_{p}^{2}+\delta}}\}}.$
Moreover, the new metric Eq. 2 indeed improves the condition number of the Riemannian Hessian, i.e., $\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))\leq\kappa_{\mathrm{e}}(\operatorname{Hess}_{\mathrm{e}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$.
@@ 5.2 RGD and RCG for truncated singular value decomposition
Let $\mathcal{M}$ be endowed with the Riemannian metric Eq. 2. We apply the Riemannian gradient descent (Algorithm 1) and Riemannian conjugate gradient (Algorithm 2) methods to solve the SVD problem Eq. 1 in Algorithms 5 and 6. Note that the retraction mapping is based on the QR decomposition, i.e., $\operatorname{R}_{(\mathbf{U},\mathbf{V})}(\eta):=(\operatorname{qf}(\mathbf{U}+\eta_{1}),\operatorname{qf}(\mathbf{V}+\eta_{2}))$ for $\eta\in\operatorname{T}_{(\mathbf{U},\mathbf{V})}\!\mathcal{M}$. The vector transport in Algorithm 6 is defined by the projection operator Eq. 3.
@@ 5.3 Numerical validation
We compare the performance of Algorithms 5 and 6 with RGD and RCG under the Euclidean metric in [23]. The proposed preconditioned metric Eq. 2, which has a right preconditioning effect, is denoted by “(R12)”. We set $m=1000$, $n=500$, $p=10$, and $\mathbf{N}=\operatorname{diag}(p,p-1,\dots,1)$. The matrix $\mathbf{A}$ is constructed by $\mathbf{A}=\mathbf{U}^{*}\Sigma(\mathbf{V}^{*})^{\mathsf{T}}$, where the entries of $\mathbf{U}^{*}\in\mathbb{R}^{m\times p}$ and $\mathbf{V}^{*}\in\mathbb{R}^{n\times p}$ are firstly sampled i.i.d. from the uniform distribution on $[0,1]$, and $\mathbf{U}^{*}$ and $\mathbf{V}^{*}$ are orthogonalized by QR decomposition. We set $\Sigma:=\operatorname{diag}(1,\gamma,\gamma^{2},\dots,\gamma^{p-1})$ and $\gamma=1/1.5$. The implementation of RGD and RCG is the same as Section 4.
@@ 7 Conclusions and future works
In the future, we intend to specify the proposed framework to other problems, e.g., the distributed PCA. Moreover, taking into account the block structure of product manifolds, parallel computing is capable of further accelerating the Riemannian optimization methods.
@@ Acknowledgments
$g_{\lambda,\mathbf{x}}=\langle\xi,\mathbf{B}_{\lambda}\eta\rangle\quad\text{for }\xi,\eta\in\operatorname{T}_{\mathbf{x}}\!\mathcal{M}_{\mathbf{B}},$
@@
$\mathrm{grad}_{g_{\lambda,\mathbf{x}}}f(x)=-(\mathbf{I}_{n}-\frac{\mathbf{B}_{\lambda}^{-1}\mathbf{B}\mathbf{x}\mathbf{x}^{\mathsf{T}}\mathbf{B}}{\mathbf{x}^{\mathsf{T}}\mathbf{B}\mathbf{B}_{\lambda}^{-1}\mathbf{B}\mathbf{x}})\mathbf{B}_{\lambda}^{-1}\mathbf{b}=-\mathbf{B}_{\lambda}^{-1}\mathbf{b}+\frac{\mathbf{x}^{\mathsf{T}}\mathbf{B}\mathbf{B}_{\lambda}^{-1}\mathbf{b}}{\mathbf{x}^{\mathsf{T}}\mathbf{B}\mathbf{B}_{\lambda}^{-1}\mathbf{B}\mathbf{x}}\mathbf{B}_{\lambda}^{-1}\mathbf{B}\mathbf{x},$
by following Proposition 4.2. Subsequently, the update rule for RGD under the metric $g_{\lambda,\mathbf{x}}$ is $\mathbf{x}^{(t+1)}={\bar{\mathbf{x}}^{(t)}}/{\|\bar{\mathbf{x}}^{(t)}\|_{\mathbf{B}}}$ with $\bar{\mathbf{x}}^{(t)}=\mathbf{x}^{(t)}-s^{(t)}\mathrm{grad}_{g_{\lambda,\mathbf{x}}}f(\mathbf{x}^{(t)})$, where we adopt the polar retraction [26, (3.3)]. The Riemannian Hessian of $f$ at $\mathbf{x}^{*}$ along $\eta\in\operatorname{T}_{\mathbf{x}^{*}}\!\mathcal{M}_{\mathbf{B}}$ can be computed by
$\operatorname{Hess}_{g_{\lambda,\mathbf{x}}}\!f(\mathbf{x}^{*})[\eta]$ $=\Pi_{g_{\lambda,\mathbf{x}},\mathbf{x}^{*}}(\mathrm{D}\mathrm{grad}_{g_{\lambda,\mathbf{x}}}f(\mathbf{x}^{*})[\eta])=\Pi_{g_{\lambda,\mathbf{x}},\mathbf{x}^{*}}(\|\mathbf{B}^{-1}\mathbf{b}\|_{\mathbf{B}}{\mathbf{B}_{\lambda}}^{-1}\mathbf{B}\eta),$
since $\mathrm{grad}_{g_{\lambda,\mathbf{x}}}f(\mathbf{x}^{*})=\Pi_{g_{\lambda,\mathbf{x}},\mathbf{x}^{*}}(-\mathbf{B}_{\lambda}^{-1}\mathbf{b})=0$. The Rayleigh quotient Eq. 4 is given by
@@
Consequently, we can compute the condition number of $\operatorname{Hess}_{g_{\lambda,\mathbf{x}}}\!f(\mathbf{x}^{*})$ in a same fashion as Proposition 4.4. Note that if $\lambda=0$, the Rayleigh quotient boils down to a constant $\|\mathbf{B}^{-1}\mathbf{b}\|_{\mathbf{B}}$ and thus $\kappa_{g_{0,\mathbf{x}}}(\operatorname{Hess}_{g_{0,\mathbf{x}}}\!f(\mathbf{x}^{*}))=1$.
@@
where $\mathbf{M}=\Sigma_{xx}^{-1/2}\Sigma_{xy}\Sigma_{yy}^{-1/2}$, $\tilde{\eta}_{1}=\Sigma_{xx}^{1/2}\eta_{1}$ and $\tilde{\eta}_{2}=\Sigma_{yy}^{1/2}\eta_{2}$. By using Eq. 4, $\mathbf{U}^{*}=\Sigma_{xx}^{-1/2}\bar{\mathbf{U}}$ and $\mathbf{V}^{*}=\Sigma_{yy}^{-1/2}\bar{\mathbf{V}}$, we can represent $\tilde{\eta}$ by
(3) $\tilde{\eta}=(\Sigma_{xx}^{\frac{1}{2}}\eta_{1},\Sigma_{yy}^{\frac{1}{2}}\eta_{2})=(\bar{\mathbf{U}}\boldsymbol{\Omega}_{1}+\bar{\mathbf{U}}_{\perp}\mathbf{K}_{1},\bar{\mathbf{V}}\boldsymbol{\Omega}_{2}+\bar{\mathbf{V}}_{\perp}\mathbf{K}_{2}),$
where $\bar{\mathbf{U}}_{\perp}\in\mathrm{St}(d_{x}-m,d_{x})$ and $\bar{\mathbf{V}}_{\perp}\in\mathrm{St}(d_{y}-m,d_{y})$ satisfy $\bar{\mathbf{U}}^{\mathsf{T}}\bar{\mathbf{U}}_{\perp}=0$, $\bar{\mathbf{V}}^{\mathsf{T}}\bar{\mathbf{V}}_{\perp}=0$ and $\mathbf{M}=[\bar{\mathbf{U}}\ \bar{\mathbf{U}}_{\perp}]\operatorname{diag}(\sigma_{1},\dots,\sigma_{r},0,\dots,0)[\bar{\mathbf{V}}\ \bar{\mathbf{V}}_{\perp}]^{\mathsf{T}}$ with $r=\mathrm{rank}(\mathbf{M})$. Taking Eq. 3 into Eq. 2, we obtain that
@@
$(\operatorname{T}_{\mathbf{U}}\!\mathcal{M}_{1})^{\perp}=\{\mathbf{U}\mathbf{S}_{1}\mathbf{M}_{1,2}^{-1}:\mathbf{S}_{1}\in\mathbb{R}^{m\times m},\ \mathbf{S}_{1}=\mathbf{S}_{1}^{\mathsf{T}}\},$

### v3

@@ 4.2 New left and right preconditioner
$\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ $=\frac{\bar{v}_{i^{*},i^{*}+1}(\delta)}{\underline{v}_{i^{*},i^{*}+1}(\delta)}=\frac{\frac{1}{2}(\mu_{i^{*}}+\mu_{i^{*}+1})(\sigma_{i^{*}}+\sigma_{i^{*}+1})}{\frac{1}{2}(\mu_{i^{*}}-\mu_{i^{*}+1})(\sigma_{i^{*}}-\sigma_{i^{*}+1})}$
$\leq\frac{\max\{\frac{1}{2}(\mu_{1}+\mu_{2})(\sigma_{1}+\sigma_{2}),\mu_{1}(\sigma_{1}+\sigma_{m+1})\}}{\min\{\underset{i,j\in[m],i\neq j}{\min}\frac{1}{2}(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j}),\mu_{m}(\sigma_{m}-\sigma_{m+1})\}}$
$=\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*})).$
If $\{i_{1}^{*},i_{2}^{*}\}\subseteq\operatorname{arg\,max}_{i\in[m]}\bar{v}_{i,i+1}(0)$ for $i_{1}^{*}<i_{2}^{*}$, it follows from the continuity of $\bar{v}_{ij}$ and $\underline{v}_{ij}$ that there exists $\bar{\delta}_{i_{1}^{*},i_{2}^{*}}>0$, such that $\bar{v}_{i_{1}^{*},i_{1}^{*}+1}(\delta)>\bar{v}_{i_{2}^{*},i_{2}^{*}+1}(\delta)$ and $\underline{v}_{i_{1}^{*},i_{1}^{*}+1}(\delta)>\underline{v}_{i_{2}^{*},i_{2}^{*}+1}(\delta)$. The rationales behind these are $\bar{v}_{i_{1}^{*},i_{1}^{*}+1}(0)=\bar{v}_{i_{2}^{*},i_{2}^{*}+1}(0)$, $\underline{v}_{i_{1}^{*},i_{1}^{*}+1}(0)=\underline{v}_{i_{2}^{*},i_{2}^{*}+1}(0)$, and the derivatives satisfy
$\bar{v}_{i_{1}^{*},i_{1}^{*}+1}^{\prime}(0)$ $=-\frac{\bar{v}_{i_{1}^{*},i_{1}^{*}+1}(0)}{2\mu_{i_{1}^{*}}\sigma_{i_{1}^{*}}\mu_{i_{1}^{*}+1}\sigma_{i_{1}^{*}+1}}>-\frac{\bar{v}_{i_{2}^{*},i_{2}^{*}+1}(0)}{2\mu_{i_{2}^{*}}\sigma_{i_{2}^{*}}\mu_{i_{2}^{*}+1}\sigma_{i_{2}^{*}+1}}=\bar{v}_{i_{2}^{*},i_{2}^{*}+1}^{\prime}(0),$
$\underline{v}_{i_{1}^{*},i_{1}^{*}+1}^{\prime}(0)$ $=-\frac{\underline{v}_{i_{1}^{*},i_{1}^{*}+1}(0)}{2\mu_{i_{1}^{*}}\sigma_{i_{1}^{*}}\mu_{i_{1}^{*}+1}\sigma_{i_{1}^{*}+1}}>-\frac{\underline{v}_{i_{2}^{*},i_{2}^{*}+1}(0)}{2\mu_{i_{2}^{*}}\sigma_{i_{2}^{*}}\mu_{i_{2}^{*}+1}\sigma_{i_{2}^{*}+1}}=\underline{v}_{i_{2}^{*},i_{2}^{*}+1}^{\prime}(0)$
if $i_{2}^{*}<m$ and
$\bar{v}_{i_{1}^{*},i_{1}^{*}+1}^{\prime}(0)$ $=-\frac{\bar{v}_{i_{1}^{*},i_{1}^{*}+1}(0)}{2\mu_{i_{1}^{*}}\sigma_{i_{1}^{*}}\mu_{i_{1}^{*}+1}\sigma_{i_{1}^{*}+1}}>-\frac{\bar{v}_{m,m+1}(0)}{2\mu_{m}^{2}\sigma_{m}^{2}}=\bar{v}_{m,m+1}^{\prime}(0),$
$\underline{v}_{i_{1}^{*},i_{1}^{*}+1}^{\prime}(0)$ $=-\frac{\underline{v}_{i_{1}^{*},i_{1}^{*}+1}(0)}{2\mu_{i_{1}^{*}}\sigma_{i_{1}^{*}}\mu_{i_{1}^{*}+1}\sigma_{i_{1}^{*}+1}}>-\frac{\underline{v}_{m,m+1}(0)}{2\mu_{m}^{2}\sigma_{m}^{2}}=\underline{v}_{m,m+1}^{\prime}(0)$
if $i_{2}^{*}=m$. Therefore, there exists $\bar{\delta}\in(0,\min\{\bar{\delta}_{i,j}:i,j\in\operatorname{arg\,max}_{i\in[m]}\bar{v}_{i,i+1}(0)\})$ and $i^{*},j^{*}\in[m]$, such that: 1) $i^{*}=\operatorname{arg\,max}_{i\in[m]}\bar{v}_{i,i+1}(\delta)$; 2) $j^{*}=\operatorname{arg\,min}_{i\in[m]}\underline{v}_{i,i+1}(\delta)$ for all $\delta\in[0,\bar{\delta})$; 3) $i^{*}<j^{*}$. Consequently, we obtain that
@@
$~{}~{}~{}~{}\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))=\frac{\bar{v}_{i^{*},i^{*}+1}(\delta)}{\underline{v}_{j^{*},j^{*}+1}(\delta)}$
$=\frac{(\mu_{i^{*}}+\mu_{i^{*}+1})(\sigma_{i^{*}}+\sigma_{i^{*}+1})}{(\mu_{j^{*}}-\mu_{j^{*}+1})(\sigma_{j^{*}}-\sigma_{j^{*}+1})}\cdot\frac{\sqrt{\mu_{j^{*}}^{2}\sigma_{j^{*}}^{2}+1}+\sqrt{\mu_{j^{*}+1}^{2}\sigma_{j^{*}+1}^{2}+1}}{\sqrt{\mu_{i^{*}}^{2}\sigma_{i^{*}}^{2}+1}+\sqrt{\mu_{i^{*}+1}^{2}\sigma_{i^{*}+1}^{2}+1}}$
$<\frac{\max\{\frac{1}{2}(\mu_{1}+\mu_{2})(\sigma_{1}+\sigma_{2}),\mu_{1}(\sigma_{1}+\sigma_{m+1})\}}{\min\{\underset{i,j\in[m],i\neq j}{\min}\frac{1}{2}(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j}),\mu_{m}(\sigma_{m}-\sigma_{m+1})\}}$
$=\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*})).$
@@ 4.3 RGD and RCG for canonical correlation analysis
Note that 1) the retraction mapping is the generalized QR factorization [22] with respect to $\Sigma_{xx}$ and $\Sigma_{yy}$, i.e.,
@@
where $\operatorname{qf}(\mathbf{X})$ refers to the $\mathbf{Q}$ factor in the QR factorization $\mathbf{Q}\mathbf{R}=\mathbf{X}$. In practice, the retraction can be efficiently computed [22] by $\operatorname{R}_{(\mathbf{U},\mathbf{V})}(\eta)=((\mathbf{U}+\eta_{1})\mathbf{R}_{1}^{-1},(\mathbf{V}+\eta_{2})\mathbf{R}_{2}^{-1})$ instead, where $\mathbf{R}_{1}^{\mathsf{T}}\mathbf{R}_{1}=(\mathbf{U}+\eta_{1})^{\mathsf{T}}\Sigma_{xx}(\mathbf{U}+\eta_{1})$ and $\mathbf{R}_{2}^{\mathsf{T}}\mathbf{R}_{2}=(\mathbf{V}+\eta_{2})^{\mathsf{T}}\Sigma_{yy}(\mathbf{V}+\eta_{2})$ are Cholesky factorization; 2) the vector transport in Algorithm 4 is defined by the projection operator in Eq. 8, i.e., $\mathcal{T}_{t\leftarrow t-1}(\eta)=\Pi_{g,(\mathbf{U}^{(t)},\mathbf{V}^{(t)})}(\eta)$ for $\eta\in\operatorname{T}_{(\mathbf{U}^{(t-1)},\mathbf{V}^{(t-1)})}\!\mathcal{M}$.
@@ 4.4 Numerical validation
5 Application to truncated singular value decomposition
@@ 5.1 A new preconditioned metric
The effect of the new metric Eq. 2 is illustrated by the following proposition, which can be proved in a same fashion by letting $\Sigma_{xx}=\mathbf{I}_{d_{x}}$, $\Sigma_{yy}=\mathbf{I}_{d_{y}}$, $\Sigma_{xy}=\mathbf{A}$, $d_{x}=m$ and $d_{y}=n$ in Propositions 4.1 and 4.4.
@@
$\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ $=\frac{\max\{\underset{i,j\in[p],i\neq j}{\max}\frac{(\mu_{i}+\mu_{j})(\sigma_{i}+\sigma_{j})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta}},\underset{i\in[p]}{\max}\frac{\mu_{i}(\sigma_{i}+\sigma_{p+1})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}}\}}{\min\{\underset{i,j\in[p],i\neq j}{\min}\frac{(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta}},\underset{i\in[p]}{\min}\frac{\mu_{i}(\sigma_{i}-\sigma_{p+1})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}}\}}.$
Moreover, the new metric Eq. 2 indeed improves the condition number of the Riemannian Hessian, i.e.,
$\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))\leq\kappa_{\mathrm{e}}(\operatorname{Hess}_{\mathrm{e}}\!f(\mathbf{U}^{*},\mathbf{V}^{*})).$
@@ 5.2 RGD and RCG for truncated singular value decomposition
Let $\mathcal{M}$ be endowed with the Riemannian metric Eq. 2. We apply the Riemannian gradient descent (Algorithm 1) and Riemannian conjugate gradient (Algorithm 2) methods to solve the SVD problem Eq. 1 in Algorithms 5 and 6. Note that the retraction mapping is based on the QR factorization, i.e., $\operatorname{R}_{(\mathbf{U},\mathbf{V})}(\eta):=(\operatorname{qf}(\mathbf{U}+\eta_{1}),\operatorname{qf}(\mathbf{V}+\eta_{2}))$ for $\eta\in\operatorname{T}_{(\mathbf{U},\mathbf{V})}\!\mathcal{M}$. The vector transport in Algorithm 6 is defined by the projection operator Eq. 3.
@@ 5.3 Numerical validation
We compare the performance of Algorithms 5 and 6 with RGD and RCG under the Euclidean metric in [23]. The proposed preconditioned metric Eq. 2, which has a right preconditioning effect, is denoted by “(R12)”. We set $m=1000$, $n=500$, $p=10$, and $\mathbf{N}=\operatorname{diag}(p,p-1,\dots,1)$. The matrix $\mathbf{A}$ is constructed by $\mathbf{A}=\mathbf{U}^{*}\Sigma(\mathbf{V}^{*})^{\mathsf{T}}$, where the entries of $\mathbf{U}^{*}\in\mathbb{R}^{m\times p}$ and $\mathbf{V}^{*}\in\mathbb{R}^{n\times p}$ are firstly sampled i.i.d. from the uniform distribution on $[0,1]$, and $\mathbf{U}^{*}$ and $\mathbf{V}^{*}$ are orthogonalized by QR factorization. We set $\Sigma:=\operatorname{diag}(1,\gamma,\gamma^{2},\dots,\gamma^{p-1})$ and $\gamma=1/1.5$. The implementation of RGD and RCG is the same as Section 4.
@@ 7 Conclusions and future works
In the future, we intend to specify the proposed framework to other problems. Moreover, taking into account the block structure of product manifolds, parallel computing is capable of further accelerating the Riemannian optimization methods.
@@ Acknowledgments
$g_{\lambda,\mathbf{x}}(\xi,\eta)=\langle\xi,\mathbf{B}_{\lambda}\eta\rangle\quad\text{for }\xi,\eta\in\operatorname{T}_{\mathbf{x}}\!\mathcal{M}_{\mathbf{B}},$
@@
$\mathrm{grad}_{g_{\lambda}}f(x)=-(\mathbf{I}_{n}-\frac{\mathbf{B}_{\lambda}^{-1}\mathbf{B}\mathbf{x}\mathbf{x}^{\mathsf{T}}\mathbf{B}}{\mathbf{x}^{\mathsf{T}}\mathbf{B}\mathbf{B}_{\lambda}^{-1}\mathbf{B}\mathbf{x}})\mathbf{B}_{\lambda}^{-1}\mathbf{b}=-\mathbf{B}_{\lambda}^{-1}\mathbf{b}+\frac{\mathbf{x}^{\mathsf{T}}\mathbf{B}\mathbf{B}_{\lambda}^{-1}\mathbf{b}}{\mathbf{x}^{\mathsf{T}}\mathbf{B}\mathbf{B}_{\lambda}^{-1}\mathbf{B}\mathbf{x}}\mathbf{B}_{\lambda}^{-1}\mathbf{B}\mathbf{x},$
by following Proposition 4.2. Subsequently, the update rule for RGD under the metric $g_{\lambda}$ is $\mathbf{x}^{(t+1)}={\bar{\mathbf{x}}^{(t)}}/{\|\bar{\mathbf{x}}^{(t)}\|_{\mathbf{B}}}$ with $\bar{\mathbf{x}}^{(t)}=\mathbf{x}^{(t)}-s^{(t)}\mathrm{grad}_{g_{\lambda}}f(\mathbf{x}^{(t)})$, where we adopt the polar retraction [26, (3.3)]. The Riemannian Hessian of $f$ at $\mathbf{x}^{*}$ along $\eta\in\operatorname{T}_{\mathbf{x}^{*}}\!\mathcal{M}_{\mathbf{B}}$ can be computed by
$\operatorname{Hess}_{g_{\lambda}}\!f(\mathbf{x}^{*})[\eta]$ $=\Pi_{g_{\lambda},\mathbf{x}^{*}}(\mathrm{D}\mathrm{grad}_{g_{\lambda}}f(\mathbf{x}^{*})[\eta])=\Pi_{g_{\lambda},\mathbf{x}^{*}}(\|\mathbf{B}^{-1}\mathbf{b}\|_{\mathbf{B}}{\mathbf{B}_{\lambda}}^{-1}\mathbf{B}\eta),$
since $\mathrm{grad}_{g_{\lambda}}f(\mathbf{x}^{*})=\Pi_{g_{\lambda},\mathbf{x}^{*}}(-\mathbf{B}_{\lambda}^{-1}\mathbf{b})=0$. The Rayleigh quotient Eq. 4 is given by
@@
Consequently, we can compute the condition number of $\operatorname{Hess}_{g_{\lambda}}\!f(\mathbf{x}^{*})$ in a same fashion as Proposition 4.4. Note that if $\lambda=0$, the Rayleigh quotient boils down to a constant $\|\mathbf{B}^{-1}\mathbf{b}\|_{\mathbf{B}}$ and thus $\kappa_{g_{0}}(\operatorname{Hess}_{g_{0}}\!f(\mathbf{x}^{*}))=1$.
@@
where $\mathbf{M}=\Sigma_{xx}^{-1/2}\Sigma_{xy}\Sigma_{yy}^{-1/2}$, $\tilde{\eta}_{1}=\Sigma_{xx}^{1/2}\eta_{1}$ and $\tilde{\eta}_{2}=\Sigma_{yy}^{1/2}\eta_{2}$. By using Eq. 4, we can represent $\tilde{\eta}$ by
$\tilde{\eta}$ $=(\Sigma_{xx}^{\frac{1}{2}}\eta_{1},\Sigma_{yy}^{\frac{1}{2}}\eta_{2})=(\Sigma_{xx}^{\frac{1}{2}}\mathbf{U}^{*}\boldsymbol{\Omega}_{1}+\Sigma_{xx}^{\frac{1}{2}}\mathbf{U}_{\Sigma_{xx}\perp}^{*}\mathbf{K}_{1},\Sigma_{yy}^{\frac{1}{2}}\mathbf{V}^{*}\boldsymbol{\Omega}_{2}+\Sigma_{yy}^{\frac{1}{2}}\mathbf{V}_{\Sigma_{yy}\perp}^{*}\mathbf{K}_{2})$
(3) $=(\bar{\mathbf{U}}\boldsymbol{\Omega}_{1}+\bar{\mathbf{U}}_{\perp}\mathbf{K}_{1},\bar{\mathbf{V}}\boldsymbol{\Omega}_{2}+\bar{\mathbf{V}}_{\perp}\mathbf{K}_{2}),$
where $\bar{\mathbf{U}}=\Sigma_{xx}^{\frac{1}{2}}\mathbf{U}^{*}\in\mathrm{St}(m,d_{x})$, $\bar{\mathbf{V}}=\Sigma_{yy}^{\frac{1}{2}}\mathbf{V}^{*}\in\mathrm{St}(m,d_{y})$, $\bar{\mathbf{U}}_{\perp}=\mathbf{U}_{\Sigma_{xx}\perp}^{*}\in\mathrm{St}(d_{x}-m,d_{x})$ and $\bar{\mathbf{V}}_{\perp}=\Sigma_{yy}^{\frac{1}{2}}\mathbf{V}_{\Sigma_{yy}\perp}^{*}\in\mathrm{St}(d_{y}-m,d_{y})$ satisfy $\bar{\mathbf{U}}^{\mathsf{T}}\bar{\mathbf{U}}_{\perp}=0$, $\bar{\mathbf{V}}^{\mathsf{T}}\bar{\mathbf{V}}_{\perp}=0$ and $\mathbf{M}=[\bar{\mathbf{U}}\ \bar{\mathbf{U}}_{\perp}]\operatorname{diag}(\sigma_{1},\dots,\sigma_{r},0,\dots,0)[\bar{\mathbf{V}}\ \bar{\mathbf{V}}_{\perp}]^{\mathsf{T}}$ with $r=\mathrm{rank}(\mathbf{M})$ by Eq. 2. Taking Eq. 3 into Eq. 2, we obtain that
@@
(1) $(\operatorname{T}_{\mathbf{U}}\!\mathcal{M}_{1})^{\perp}=\{\mathbf{U}\mathbf{S}_{1}\mathbf{M}_{1,2}^{-1}:\mathbf{S}_{1}\in\mathbb{R}^{m\times m},\ \mathbf{S}_{1}=\mathbf{S}_{1}^{\mathsf{T}}\},$